# Word problems and embedding-obstructions in cellular automata groups on groups

### v2

@@ 4 Unconditional nonembedding results
First, consider the case $A=\mathbb{Z}_{7}\times\{0,1\}^{D-1}$. In this case, we show that $\mathrm{PAut}(A)$ can be used in the proof.
For the following discussion, fix $n=2^{k}\in\mathbb{N}$ and a configuration $x\in A^{\mathbb{Z}^{D}}$. The counters of $\vec{v}$, written $C(\vec{v})$ are the $D-1$ many numbers whose binary representations of length $k+1$ are written on the $\{0,1\}^{D-1}$-component (each cell contains one bit of each number), in the consecutive positions $(\vec{v},\vec{v}+e_{D},\ldots,\vec{v}+ne_{D})$. The components in this vector must all have a value in $[0,n]=[0,2^{k}]$, or otherwise the counters are not defined. A coordinate $\vec{v}$ is good if its counters $\vec{u}$ exist, and for each natural generator $e_{i}$ of $\mathbb{Z}^{D}$ except $e_{D}$, the counters of $\vec{v}+e_{i}$ exist and are $\vec{u}+e_{i}$.
We define analogs of automorphisms used in the proof of Theorem 3.1. Let $\pi$ be an even permutation of $\mathbb{Z}_{7}$, and define $\beta_{\pi}$ as the automorphism that performs $\pi$ on $\mathbb{Z}_{7}$ at $\vec{v}$ if the coordinate $\vec{v}$ is good. From Lemma A.15 (though the input is formatted slightly differently) we see that $\beta_{\pi}$ has word length polynomial in $n$ in the word norm of $\mathrm{PAut}(A)$.
For $\pi\in\mathrm{Alt}(\mathbb{Z}_{7})$ and $s$ not in the support of $\pi$, let $\phi_{\pi,s,i}$ be the automorphism that performs $\pi$ on $\mathbb{Z}_{7}$ at $\vec{v}$ if the $\mathbb{Z}_{7}$-value at $\vec{v}+e_{i}$ is $s$. Analogously to Equation 1 in the proof of Theorem 3.1, we see $\phi_{\pi,s,i}\in\mathrm{PAut}(A)$ with constant word norm (instead of the borrowed trit in the $\{0,1,2\}$-component, one can use one of the bits in the $\{0,1\}^{D-1}$-component as a borrowed bit that returns to its original value after the operation). Let $\gamma_{\pi,\ell}$ be the automorphism that permutes $\mathbb{Z}_{7}$ at $\vec{v}$ by $\pi$ if and only if the sum of the counters at $\vec{v}$ is exactly $\ell$. This has polynomial word norm by Lemma A.16 (again, understanding that inputs are formatted slightly differently).
Define $S_{i}=\{\vec{v}\;|\;\vec{v}\geq 0\wedge\vec{v}_{D}=0\wedge\sum\vec{v}=i\}$, a discrete analog of the standard simplex, and define $T_{i}=\bigcup_{j\leq i}S_{j}$. Note that $|T_{i}|=\sum_{j=0}^{i}\geq C^{\prime}i^{D-1}$ for some global constant $C^{\prime}$. Note also that $\gamma_{\pi,\ell}$ acts nontrivially precisely when the counter value is $\vec{v}$ such that $\vec{v}^{\prime}\in S_{\ell}$, where $\vec{v}^{\prime}_{i}=\vec{v}_{i}$ for $1\leq i<D$ and $\vec{v}^{\prime}_{D}=0$. In what follows, we often identify vectors of dimension $D-1$ with vectors of dimension $D$ with 00 in the last coordinate.
We now perform dyadically accelerated funneling along the sets $T_{i}$, analogously to the proof of Theorem 3.1. For $i\in[0,k]$ and $j\in[0,2^{k-i}-1]$, let $I=[j2^{i},(j+1)2^{i}]\subset[0,n]$, let $s\in\mathbb{Z}_{7}$ and $\pi\in\mathrm{Alt}(\mathbb{Z}_{7})$ a $3$-rotation. Note that $|I|\geq 2$ always.
We define an automorphism $\psi_{j,i,s,\pi}$ as follows for $\pi$ a $3$-rotation: The $\{0,1\}^{D-1}$-components of cells never change their values. On the first track, at $\vec{v}$, the value changes by an application of $\pi$ if and only if the following hold:
  • the counters $C(\vec{v})$ exists and $\sum C(\vec{v})=j2^{i}$ i.e. $C(\vec{v})\in S_{j2^{i}}$;
  • for all $\ell\in[0,2^{i}-1]$ and any $\vec{u}\in S_{\ell}$, the node $\vec{v}+\vec{u}$ is good;
  • for all $\ell\in[1,2^{i}-1]$ and any $\vec{u}\in S_{\ell}$, the $\mathbb{Z}_{7}$-value at $\vec{v}+\vec{u}$ is 00;
  • for all $\vec{u}\in S_{2^{i}}$, the $\mathbb{Z}_{7}$-value at $\vec{v}+\vec{u}$ is $s$.
Note that by the first two items, the counter value $C(\vec{v}+\vec{u})=j2^{i}+\ell$ for all $\ell\in[0,2^{i}]$ and any $\vec{u}\in S_{\ell}$.
Suppose first that $s$ is not in the support of $\pi$. Then for $i=0$ and $j\in[0,2^{k-i}-1]$ we can build the above automorphisms from the automorphisms $\beta_{\pi}$ and $\phi_{\pi_{1},s,1}$ described above: by letting $\pi=[\pi_{0},\pi_{1},\ldots,\pi_{D+1}]$ we have
$\psi_{j,0,s,\pi}=[\beta_{\pi_{0}},\phi_{\pi_{1},s,1},\ldots,\phi_{\pi_{D},s,D-1},\gamma_{\pi_{D+1},j}].$
The norm is polynomial in $k$ (though there is a multiplicative constant exponential in $D$).
The idea for the proof of correctness is that we can interpret the maps
$\beta_{\pi_{0}},\phi_{\pi_{1},s,1},\ldots,\phi_{\pi_{D},s,D-1},\gamma_{\pi_{D+1},j}$
as controlled permutations at a given node $\vec{v}$ checking the various requirements for $\psi_{j,0,s,\pi}$ applying, and as in Lemma A.9 the effect of the iterated commutator is to perform the iterated commutator of the $\pi_{i}$ if all these conditions hold, and otherwise do nothing. The detailed proof is exactly analogous to the corresponding claim in Theorem 3.1, but using $T_{j}$ in place of the tree $\{a,b\}^{\leq 2^{j}}$ (unlike in that proof, the shifts of $T_{j}$ appearing in the proof have nontrivial intersections, but this does not affect the logic in any way).
[…]
still has norm polynomial in $k$. Again, the detailed proof is as in Theorem 3.1. Let $N_{0}$ be an upper bound on norms of the maps $\psi_{j,0,s,\pi}$, which we have concluded is polynomial in $k$, i.e. it is $O(\log(n)^{t})$ for some $t$.
We perform induction on $i=0,\ldots,k$ to get the maps $\psi_{j,i,s,\pi}$. We have
$[j2^{i},(j+1)2^{i}]=[2j2^{i-1},(2j+1)2^{i-1}]\cup[(2j+1)2^{i-1},2(j+1)2^{i-1}]$
and the two intervals have intersection $\{(2j+1)2^{i-1}\}$. It is easy to check (see again the proof of the corresponding claim in the proof of Theorem 3.1) that
[…]
where $\pi=[\pi_{1},\pi_{2}]$, and $\pi_{3}$ is any permutation taking 00 to $1$.
Writing $N_{i}$ for the maximal word norm of such an element $\psi_{j,i,s,\pi}$, the formula shows $N_{i}\leq 8N_{i-1}$, so $N_{k}\leq 8^{k}N_{0}=O(n^{3}\log(n)^{t})$ for some $t$.
The automorphism $\psi_{0,n,\pi,0}$ can now fully described as follows: Say $\vec{v}$ has a marker (and $\vec{v}$ is the root of the marker) if for all $\vec{u}\in T_{n}$, the counter values in $\vec{v}+\vec{u}$ are $\vec{u}$. Note that this provides a full description of the contents of $\vec{v}+\vec{u}+\vec{w}$ for $\vec{w}\in[0,k]e_{D},\vec{u}\in T_{n}$). Enumerate the $M=\Omega(n^{D-1})$ vectors in $T_{n}$ as $\vec{t}_{1},\ldots,\vec{t}_{M}$, starting with $\vec{t}_{1}=\vec{0}$. Then $\psi_{0,n,\pi,0}$ is the automorphism that, whenever $\vec{v}$ has a marker, performs the $3$-rotation
$((a,0,\ldots 0);\;(b,0,\ldots 0);\;(c,0,\ldots 0))$
on the $\mathbb{Z}_{7}$-symbols read in positions $\vec{t}_{1},\ldots,\vec{t}_{M}$, where $\pi=(a\;b\;c)$.
Thus, $\psi_{0,n,\pi,0}$ is an element of the simple group of all even permutations of such vectors (performed on the $\mathbb{Z}_{7}$-components of cells $\vec{v}+T_{n}$ when $\vec{v}$ has a marker). For this to be well-defined, it is crucial that the automorphisms never change the positions of markers, and that the $\mathbb{Z}_{7}$-symbols being affected by the permutations cannot belong to markers at two different roots $\vec{v},\vec{v}^{\prime}$ such that both have a marker. But this is clear because we write the counters explicitly, never change the values of counters, and they describe the position of each cell in the corresponding marker $\vec{v}+T_{n}$, thus identifying the root $\vec{v}$.
We conclude that in $\mathrm{PAut}_{\mathbb{Z}^{D}}(A)$, there is an element of norm at most $C^{\prime\prime}n^{3}\log(n)^{t}$ for some global constants $C^{\prime\prime},t$, which belongs to a simple group $H$ of order $27^{M}!/$ 2 where $M=C^{\prime}n^{D-1}$ for a global constant $C^{\prime}$.
Suppose now for a contradiction that $\mathrm{PAut}_{\mathbb{Z}^{D}}(A)$ embeds in $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$ for some alphabet $B$. By Lemma 4.1, there exists a constant $C$ such that if $H\leq\mathrm{Aut}(A^{\mathbb{Z}^{d}})$ is a simple group containing a nontrivial element of norm at most $N=C^{\prime\prime}n^{3}\log(n)^{t}$, we have $|H|\leq C^{N^{d}}!$. But a short calculation shows that
$|H|=7^{C^{\prime}n^{D-1}}!/2>C^{(C^{\prime\prime}n^{3}\log(n)^{t})^{d}}!$
for large $n$ when $D-1>3d$. Namely, these are the respective degrees of double exponentials appearing in the formulas. This is a contradiction, and proves that no such embedding can exist.
To get the result for general alphabet $A^{\prime}\ni 0,1$, we can use words of the form $0110w$ 0 with $w$ not containing $11$ as a subword, to encode symbols of of $A^{2}$ where $A$ is the above alphabet (the square is for technical reasons, and the coding can be arbitrary), write $W$ for the set of these words, with $|W|=|A|^{2}$. Note that this set of words is mutually unbordered. The idea is that we use only the first encoded $A$-track in the argument, and the counters will be again written in consecutive positions in dimension $D$ as in the above proof, but now using $r(D+1)$ positions instead of $D+1$ as before. I.e. we use a constant number of such words $0110w$ 0 with $w\in\{0,1\}^{r-4}$ for fixed $r$ to encode the counters.
We replace the group $\mathrm{PAut}$ as follows: for symbol permutations, permute the symbols interpreted inside marker; as for partial shifts in direction $e_{i}$, we use the conveyor belt trick (see e.g. [28] for a more detailed discussion): if $i<D$, then in every chain of markers in direction $e_{i}$ (in the same phase, i.e. with the $11$-subwords in consecutive positions in direction $e_{i}\mathbb{Z}$), in each marker we interpret two symbols $s,s^{\prime}$ on the two tracks $A^{2}$, and shift the first symbol $s$ in direction $e_{i}$, and $s^{\prime}$ in direction $-e_{i}$. At the ends of such chains, we have a unique natural way to continue this to a bijection, by moving symbols between tracks, so they move as if on a conveyor belt. If $i=D$, we do the same, but now we look at progressions of consecutive markers with difference $e_{D}D$.
Now, the above argument goes through essentially verbatim, when we assume that all the controlled rotations are performed only on symbols on the first $A^{2}$-track (partial shifts may move $\mathbb{Z}_{7}$-symbols to the other track if they reach the end of the conveyor belts, but in this case, all the commutators will cancel in the end, and this $\mathbb{Z}_{7}$-symbol is not affected by the final automorphism). ∎
The case of groups containing free submonoids is similar, and we outline how to obtain it from the proof of Theorem 3.1.
Theorem 4.2.
Let $G$ be a f.g. group containing a free submonoid on two generators, and let $A$ be any alphabet. Then $\mathrm{Aut}(A^{G})$ has an f.g. subgroup that does not embed into $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$ for any alphabet $B$ and $d\in\mathbb{N}$.
Proof.
Assume $A$ has many prime divisors. In the case of a general alphabet, the marker argument is similar as in the previous theorem.
We explain only informally how the dyadically accelerated funneling argument is applied, i.e. the modifications made in the definition of the $\psi$-maps (and implicitly, the notion of goodness) in the proof of Theorem 3.1. First, the Turing machines $T_{1}$ and $T_{2}$ can now simply behave trivially, and accept all words. (These Turing machines play no role in this argument, and their states and tape symbols need not be included in the alphabets.)
Then $\psi$ simply rotates the $\mathbb{Z}_{7}$-symbol at the root of a tree where the nodes are all good and the other $\mathbb{Z}_{7}$-coordinates contain 00, and has polynomial norm in $n$. In particular, there is only one possible projection of such a tree on the non-$\mathbb{Z}_{7}$-coordinates, and furthermore, if we have such trees at $h,h^{\prime}$ with $h\neq h^{\prime}$, then $h\{a,b\}^{\leq n}\cap h^{\prime}\{a,b\}^{\leq n}=\emptyset$ (since the element of $\{a,b\}^{*}$ is visible in the index).
Thus, there is a well-defined group $H$ of even permutations on the $\mathbb{Z}_{7}$ track under such trees. Since the tree itself has $12^{n+1}-$ 1 nodes, the number of configurations is $7^{2^{n+1}-1}$, and thus we have a simple group of cardinality $|H|=7^{2^{n+1}-1}!/2$ with an element of polynomial norm. By Lemma 4.1, and a short calculation, we cannot have an embedding of $H$ into any group of the form $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$. ∎
Remark 4.1.
One could generalize results of this section to other subshifts and groups, with a bit of additional work, although we do not know whether this would lead to any striking results. There are some subtleties, e.g. in $\mathbb{Z}^{d}$-SFTs, it is surely possible that periodic points are dense, but patterns of shape $[0,n]^{D}$ do not fit into boxes of size computable in $n$. The proof of denseness of periodic points in FEP subshifts in [25] shows that at least on $\mathbb{Z}^{d}$, an analog of Lemma 4.1 does hold. In the case of general residually finite groups, even on full shifts it may not be possible to find efficient quotients to use in the proof. Furthermore, our proofs here also use rigid cones in the case of $\mathbb{Z}^{d}$ and embedded semigroups in the case of other groups. It should be straightforward to replace these arguments by the use of abstractly embedded trees, as in the following sections.
@@ 4.1 Dropping the centers
The following is due to Hochman [19]. The one-dimensional result is due to Ryan [27].
@@
Suppose that $D>3d+1$, $\mathcal{G}_{d}=\mathrm{Aut}(A^{\mathbb{Z}^{d}}),\mathcal{G}_{D}=\mathrm{Aut}(B^{\mathbb{Z}^{D}})$ for any nontrivial alphabets $A,B$. Then $\mathcal{G}_{D}$ has a finitely-generated subgroup such that $\mathcal{G}_{D}/K$ does not embed in $\mathcal{G}_{d}/K^{\prime}$ for any central subgroups $K\leq\mathcal{G}_{D},K^{\prime}\leq\mathcal{G}_{d}$.
@@
A we showed, there is a subgroup $H$ of $\mathcal{G}_{D}$ which is nonabelian simple of order $27^{\Theta(n^{D-1})}!/$ 2, and contains an element of word norm $n$. Now, $H$ must have trivial intersection with $K$, since otherwise $K\cap H$ is a nontrivial simple subgroup of $H$. Thus by the second isomorphism theorem, $HK/K\cong H$ as a subgroup of $\mathcal{G}/K$, and we conclude that $\mathcal{G}/K$ still has a simple subgroup of order $27^{\Theta(n^{D-1})}!/$ 2 and with an element of word norm $n$.
@@
Assume now that $K^{\prime}$ is of this form, and let $L=\langle e_{d^{\prime}+1},\ldots,e_{d^{\prime\prime}}\rangle$. Consider a finitely-generated subgroup $H^{\prime}$ of $\mathcal{G}_{d}/K^{\prime}$. We may take representatives for the generators $\mathcal{S}^{\prime}\subset\mathcal{G}_{d}$, and suppose $r$ is a bound on their radii. Suppose that some nontrivial element $h\in H^{\prime}$ of this group has word norm $n$ and belongs to a nonabelian simple group of order $27^{\Theta(n^{D-1})}!/$ 2. Let $h^{\prime}$ be the corresponding representative in $\mathcal{G}_{d}$ (of the same word norm). Since $h$ is nontrivial, $h^{\prime}$ is not contained in $K^{\prime}$. Note that the radius of $h^{\prime}$ is at most $rn$.
@@
We now claim that we can find a point with periods at most $10rn$ in each direction, such that $h^{\prime}$ does not act trivially on the $K^{\prime}$-orbits of periodic points. We observe that $\mathcal{G}_{d}/K^{\prime}$ has a well-defined action on the $K^{\prime}$ orbits of such points, and thus so does $H^{\prime}$. Since $H^{\prime}$ is simple, as in the proof of Lemma 4.1 we see that $H^{\prime}$ must act faithfully on the set of $K^{\prime}$-orbits, and we conclude that again $|H^{\prime}|\leq C^{n^{d}}!$ for some constant $C$ (which is only a function of $r$ and the alphabet. Note here that the constant $10$ also disappears into $C$, and that the number of $K^{\prime}$-orbits is smaller than the number of periodic points, which only gives a better upper bound for $H$.
@@
There are two cases to consider for this claim. Suppose $B\ni 0,1$. First, suppose $h^{\prime}$ is a shift by an element not in $K^{\prime}$, then it is a shift by a vector of radius at most $rn$ from outside $K^{\prime}$. We can easily find point with periods $(2rn+1)$ where $h^{\prime}$ it does not act by a $K^{\prime}$-shift, for example we can pick the point $x\in\{0,1\}^{\mathbb{Z}^{d}}$ with periods $2(rn+1)\mathbb{Z}^{d}$, where in the fundamental domain $[0,2rn]^{d}$ we have $x_{\vec{v}}=1\iff\vec{v}\in K$.
@@
Suppose then that $h^{\prime}$ is not a shift map. Let first $y\in\{0,1\}^{\mathbb{Z}^{d}}$ be the characteristic function of $\{\vec{0}\}$, i.e. $y_{\vec{v}}=1\iff\vec{v}=\vec{0}$. Then we have two subcases. First, if $h^{\prime}(y)\notin K^{\prime}\cdot y$, then it is easy to see that we can pick the point $x\in\{0,1\}^{\mathbb{Z}^{d}}$ with periods $2(rn+1)\mathbb{Z}^{d}$, where in the fundamental domain $[0,2rn]^{d}$ we have $x_{\vec{v}}=1\iff\vec{v}=\vec{0}$, and then $x$ is not mapped into its $K^{\prime}$-shift orbit.

### v3

@@ 4 Nonembedding results
By Lemma 4.5, $\mathrm{Aut}(A^{\mathbb{Z}^{D}})$ has a finitely-generated subgroup with simplistic growth at most $e^{n^{D}/\mathrm{poly}\log(n)}!$. If $\mathrm{Aut}(A^{\mathbb{Z}^{D}})\leq\mathrm{Aut}(B^{\mathbb{Z}^{d}})$, then this finitely-generated subgroup would have simplistic growth at least $e^{n^{d}}!$ by Lemma 4.3, a contradiction since
$e^{n^{D}/\mathrm{poly}\log(n)}!\not\lesssim e^{n^{d}}!.$
@@ 4.1 Dropping the centers
The following is due to Hochman [21] in general dimension. The one-dimensional result is due to Ryan [27].
@@
Suppose that $D>d$, $\mathcal{G}_{d}=\mathrm{Aut}(A^{\mathbb{Z}^{d}}),\mathcal{G}_{D}=\mathrm{Aut}(B^{\mathbb{Z}^{D}})$ for any nontrivial alphabets $A,B$. Then $\mathcal{G}_{D}$ has a finitely-generated subgroup such that $\mathcal{G}_{D}/K$ does not embed in $\mathcal{G}_{d}/K^{\prime}$ for any central subgroups $K\leq\mathcal{G}_{D},K^{\prime}\leq\mathcal{G}_{d}$.
@@
A we showed, there is a subgroup $\mathcal{H}$ of $\mathcal{G}_{D}$ which is nonabelian simple of order $e^{\Omega(n^{D})}!$, and contains an element of word norm $n^{\prime}=O(n\mathrm{poly}(\log(n))$. Now, $\mathcal{H}$ must have trivial intersection with $K$, [^3] Here note that since $\mathbb{Z}^{d}$ is abelian, shift maps are automorphisms, and we identify them with a subgroup of $\mathrm{Aut}$. since otherwise $K\cap\mathcal{H}$ is a nontrivial simple subgroup of $\mathcal{H}$. Thus by the second isomorphism theorem, $\mathcal{H}K/K\cong\mathcal{H}$ as a subgroup of $\mathcal{G}/K$, and we conclude that $\mathcal{G}/K$ still has a simple subgroup of order $e^{\Omega(n^{D})}!$ and with an element of word norm $n^{\prime}$.
@@
Assume now that $K^{\prime}$ is of this form, and let $L=\langle e_{d^{\prime}+1},\ldots,e_{d^{\prime\prime}}\rangle$. Consider a finitely-generated subgroup $H^{\prime}$ of $\mathcal{G}_{d}/K^{\prime}$. We may take representatives for the generators $\mathcal{S}^{\prime}\subset\mathcal{G}_{d}$, and suppose $r$ is a bound on their radii. Suppose that some nontrivial element $h\in\mathcal{H}^{\prime}$ of this group has word norm $n^{\prime}$ and belongs to a nonabelian simple group of order $e^{\Omega(n^{D})}!$. Let $h^{\prime}$ be the corresponding representative in $\mathcal{G}_{d}$ (of the same word norm). Since $h$ is nontrivial, $h^{\prime}$ is not contained in $K^{\prime}$. Note that the radius of $h^{\prime}$ is at most $rn^{\prime}$.
@@
We observe that $\mathcal{G}_{d}/K^{\prime}$ has a well-defined action on the set of $K^{\prime}$ orbits of points, and thus so does $\mathcal{H}^{\prime}$. We claim that we can find a point with periods at most $10rn^{\prime}$ in each direction, such that $h^{\prime}$ does not act trivially on its $K^{\prime}$-orbit. Since $\mathcal{H}^{\prime}$ is simple, as in the proof of Lemma 4.1 we see that $H^{\prime}$ must act faithfully on the set of $K^{\prime}$-orbits with periods at most $10rn^{\prime}$, and we conclude that again $|\mathcal{H}^{\prime}|\leq e^{O((10rn^{\prime})^{d})}!=e^{O((n^{\prime})^{d})}!$ (the invisible constant is only a function of $r$ and the alphabet $B$). This concludes the proof by again observing that $e^{O((n^{\prime})^{d})}!$ is far smaller than $e^{\Omega(n^{D})}!$.
@@
We just need to prove the claim about the action on periodic points. Suppose $B\ni 0,1$. There are two cases to consider. First, suppose $h^{\prime}$ is a shift by an element not in $K^{\prime}$, then it is a shift by a vector of radius at most $rn$ from outside $K^{\prime}$. We can easily find a point with periods $(2rn^{\prime}+1)$ where $h^{\prime}$ does not act by a $K^{\prime}$-shift, for example we can pick the point $x\in\{0,1\}^{\mathbb{Z}^{d}}$ with periods $2(rn^{\prime}+1)\mathbb{Z}^{d}$, where in the fundamental domain $[0,2rn^{\prime}]^{d}$ we have $x_{\vec{v}}=1\iff\vec{v}=\vec{0}$, and then $x$ is not mapped into its $K^{\prime}$-shift orbit.
@@
Suppose then that $h^{\prime}$ is not a shift map. Let first $y\in\{0,1\}^{\mathbb{Z}^{d}}$ be the characteristic function of $\{\vec{0}\}$, i.e. $y_{\vec{v}}=1\iff\vec{v}=\vec{0}$. Then we have two subcases. First, if $h^{\prime}(y)\notin K^{\prime}\cdot y$, then we can pick the same point $x\in\{0,1\}^{\mathbb{Z}^{d}}$ as above.